1/3 + 5/12 + 4/5 =
help me please still do not no how to add/subtract fractions...Do all the bottom numbers have to be the same number? Then all the top numbers have to be mutliplied by the same as the bottom number???

Answer by edjones(7311)

All the bottom numbers (denominators) have to be the same (common denominator).
What number do 3, 5, and 12 go into? 60 is one.
All we have to do now is divide the denominator of each original number into the common denominator and multiply the result times the original numerator to get the new numerator.
For 1/3 divide 3 into 60 = 20 and 20 times the numerator of 1 = 20
The 1st new fraction is 20/60
Next 12 goes into 60 5 times. 5 times 5 = 25. 2nd fraction is 25/60
4/5 = 48/60
20/60 + 25/60 + 48/60 = 93/60 = 31/20 or 1 11/20 Both are correct.
Ed